OpenStudy (anonymous):

A 7-foot ladder is leaned against a building in such a way that the bottom of the ladder is 4 feet from the base of the wall. Find the angle formed between the ladder and the ground. A.30 B.35 C.55 D.60

OpenStudy (anonymous):

\[\theta =\frac{180 \cos ^{-1}\left(\frac{4}{7}\right)}{\pi }=55.1501 \text{ degrees} \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Cos(x)=4/7 Cos(x)=.57 Take inverse of cos(.57) 55.1501 degrees

OpenStudy (anonymous):

The Mathematica program was used for the calculation. Mathematica computes using radians for angle measurement.

OpenStudy (anonymous):

Your welcome (:

OpenStudy (anonymous):

so whats the awnser lol